# Pairing in Fission: Studies of Static Approach and Collective Inertias

### v1

@@ Acknowledgements.
This research is part of the project No. 2021/43/P/ST2/03036 co-funded by the National Science Centre and the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement no. 945339. The work of LMR and SAG is supported by Spanish Agencia Estatal de Investigacion (AEI) of the Ministry of Science and Innovation (MCIN) under Grant No. PID2021-127890NB-I00. SG acknowledges support funded by MCIN/AEI/10.13039/501100011033 and the “European Union NextGenerationEU/PRTR” under grant agreement No. RYC2021-031880-I.
@@ Appendix A On the sign of occupation factor
Figure 9: The relation between $\Delta$ and $\langle\Delta N^{2}\rangle$ for 262Rf.
@@
We point out that BCS wave function differing in the sign of the $v_{k}$ do no longer have overlap one. The overlap between such wave functions is given by the product extended to all quantum numbers in the basis of the factors $u_{k}u^{\prime}_{k}+v_{k}v^{\prime}_{k}$. For those quantum numbers where the occupancy $v^{\prime}_{k}$ has the same sign as $v_{k}$, the factor is $u_{k}u_{k}+v_{k}v_{k}=1$. However, for those with opposite sign, the factor becomes $u_{k}u_{k}-v_{k}v_{k}\leq 1$, and therefore they contribute to decrease the overlap.

### v2

@@ Acknowledgements.
This research is part of the project No. 2021/43/P/ST2/03036 co-funded by the National Science Centre and the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement no. 945339. The work of LMR and SAG is supported by Spanish Agencia Estatal de Investigacion (AEI) of the Ministry of Science and Innovation (MCIN) under Grant No. PID2021-127890NB-I00. SG acknowledges support funded by MCIN/AEI/10.13039/501100011033 and the “European Union NextGenerationEU/PRTR” under grant agreement No. RYC2021-031880-I.
@@ Appendix A On the sign of occupation factor
Figure 9: The relation between $\Delta$ and $\langle\Delta N^{2}\rangle$ for 262Rf. The red circles correspond to the $\Delta$ values obtained in the calculations with the constrain on given $\langle\Delta N^{2}\rangle$ and the constrain on the various $Q_{20}$ within the entire examined range (from 0 to 100 b).
@@
We point out that BCS wave function differing in the sign of the $v_{k}$ no longer have overlap equals one. The overlap between such wave functions is given by the product extended to all quantum numbers in the basis of the factors $u_{k}u^{\prime}_{k}+v_{k}v^{\prime}_{k}$. For those quantum numbers where the occupancy $v^{\prime}_{k}$ has the same sign as $v_{k}$, the factor is $u_{k}u_{k}+v_{k}v_{k}=1$. However, for those with opposite sign, the factor becomes $u_{k}u_{k}-v_{k}v_{k}\leq 1$, and therefore they contribute to decreasing the overlap.